# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 8.2 SO($N_{c}$) gauge Higgs theory with scalar fields in the fundamental representation
The NAHFT with gauge group SO($N_{c}$) and multiflavor scalar fields in the fundamental representation is obtained by considering a real scalar field $\Phi^{af}$ (with $a=1,...,N_{c}$ and $f=1,...,N_{f}$) and a gauge field ${\cal A}_{\mu}\equiv\sum_{c}A_{\mu}^{c}T^{c}$, where the matrices $T^{k}$ are the real anti-symmetric generators of the SO($N_{c}$) algebra. The most general renormalizable Lagrangian is [376, 273, 155]
${\cal L}_{\rm SO}={1\over 2g^{2}}{\rm Tr}\,{\cal F}_{\mu\nu}^{2}+{1\over 2}\sum_{af\mu}\Big{(}\partial_{\mu}\Phi^{af}+\sum_{b}{\cal A}_{\mu}^{ab}\Phi^{bf}\Big{)}^{2}+V(\Phi)\,,$ (94)
@@
where ${\cal F}_{\mu\nu}$ is the non-Abelian field strength associated with the gauge field ${\cal A}_{\mu}$. For any value of $N_{c}$ and $N_{f}$, the Lagrangian (94) is invariant under the local SO($N_{c}$) gauge transformations,
@@
The RG flow of the theory does not have stable charge FPs with $\alpha>0$, unless $N_{f}>N_{f}^{*}(d)$, where $N_{f}^{*}(d)$ is very large in four dimensions. Indeed, we find $N^{*}_{f}(4)=210.5$ for $N_{c}=2$, and $N^{*}_{f}(4)\approx 443.8$, for $N_{c}=3$. The stable FPs always lie in the region $u,v>0$. More precisely, one finds $v^{*}\approx 6\varepsilon/N_{f}$, $u^{*}\approx 6\varepsilon/N_{f}$, and $\alpha^{*}=12\varepsilon/N_{f}$ for large values of $N_{f}$. Note that the existence of a critical theory for large values of $N_{f}$ is also confirmed by the analysis of the model using the large-$N_{f}$ expansion [273].
The results reviewed above show that continuous transitions with an effective field theory description are only possible for a large number of components. Again, we remind the reader that this result does not exclude the presence of other types of continuous transitions (see Sec. 3.1) for small values of $N_{f}$.
@@ 8.3 SU($N_{c}$) gauge Higgs theory with scalar fields in the adjoint representation
To define the model we consider real scalar fields $\Phi^{af}$ with $a=1,...,N_{c}^{2}-1$ (color index) and $f=1,...,N_{f}$ (flavor index), and real gauge fields ${\cal A}_{\mu}\equiv i\sum_{c}A_{\mu}^{c}T_{A}^{c}$, where $T_{A}^{abc}=-if^{abc}$ are the Hermitian SU($N_{c}$) generators in the adjoint representation ($f^{abc}$ are the structure constants of the SU($N_{c}$) group). The most general renormalizable Lagrangian is
${\cal L}_{\rm SUadj}={1\over 2g^{2}}{\rm Tr}\,{\cal F}_{\mu\nu}^{2}+\sum_{af\mu}(\partial_{\mu}\Phi^{af}+\sum_{b}{\cal A}_{\mu}^{ab}\Phi^{bf})^{2}+V(\Phi),$ (97)
@@
where ${\cal F}_{\mu\nu}$ is the non-Abelian field strength associated with the gauge field ${\cal A}_{\mu}$. The model is invariant under left SU($N_{c}$) local transformations and under right O($N_{f}$) global transformations.
@@ 9 Lattice non-Abelian Higgs models
Like LAH theories, see Secs. 5 and 6, one of the main issue is the identification of the effective field-theoretical description of the phase transitions present in LNAH models. Their phase diagrams exhibit transitions of various types, as discussed in Sec. 3.1. Indeed, these models show transitions that admit a LGW description in terms of a gauge-invariant order-parameter field, which is appropriate when gauge modes are not critical, and GFT transitions with critical gauge modes, associated with the stable charged FPs of the NAHFTs discussed in Sec. 8. Moreover, when scalar fields transform in the adjoint representation of the SU($N_{c}$) gauge group, also topological transitions appear, which are not associated with the breaking of the global symmetry. Note that, to obtain a nonperturbative definition of the NAHFT as the continuum limit of a lattice model, it is crucial that the corresponding LNAH models, with analogous local and global symmetries, undergo charged continuous transitions where both scalar and gauge modes develop critical behaviors.
@@ 9.1.1 The lattice Hamiltonian
The Hamiltonian $H$ is the sum of the scalar-field kinetic term $H_{K}$, of the local scalar potential $H_{V}$, and of the pure-gauge Hamiltonian $H_{G}$. As usual, we set the lattice spacing equal to one, so that all lengths are measured in units of the lattice spacing. The kinetic term $H_{K}$ is given by
@@
The Hamiltonian $H$ is invariant under global U($N_{f}$) transformations that only act on the scalar field, $\Phi\to\Phi W$, $W\in\hbox{U($N_{f}$)}$, and under local SU($N_{c}$) transformations
$\Phi_{\bm{x}}\to V_{\bm{x}}\Phi_{\bm{x}}\qquad U_{{\bm{x}},\mu}\to V_{\bm{x}}U_{{\bm{x}},\mu}V_{{\bm{x}}+\hat{\mu}}^{\dagger},\qquad V_{\bm{x}}\in{\rm SU}(N_{c}).$ (103)
The lattice model (99) can be simplified by considering the fixed-length limit of the scalar field, i.e., ${\rm Tr}\,\Phi_{\bm{x}}^{\dagger}\Phi_{\bm{x}}^{\phantom{\dagger}}=1$, which can be formally obtained by considering the limit $u\to\infty$ keeping the ratio $r/u=-1$ fixed. In the unit-length limit the lattice potential (101) reduces to
@@
The model in this unit-length limit is generally expected to have the same features as models with generic $r$ and $u$.
@@
where $\epsilon^{ab}$ is the completely antisymmetric tensor in the color space, with $\epsilon^{12}=1$. Note that, if $N_{f}>N_{c}=2$, the composite operator $Y_{\bm{x}}^{fg}$ transforms nontrivially under the SU($N_{f}$) group and hence the U(1) symmetry can be spontaneously broken only in phases characterized by a broken SU($N_{f}$) symmetry. Instead, for $N_{f}=N_{c}=2$, the operator $Y_{\bm{x}}^{fg}$ can be rewritten as $Y_{\bm{x}}^{fg}=\epsilon^{fg}\det\Phi_{\bm{x}}$, and is therefore invariant under SU(2) global transformations. If $N_{f}=N_{c}$, the U(1) symmetry can thus be broken irrespectively of the breaking of the SU($N_{f}$) symmetry. Analogous results hold for generic values of $N_{c}$, see Refs. [167, 99].
9.1.2 Low-temperature Higgs phases
We now outline [99] the main features of the phase diagram of the LNAH model (99) in the unit-length limit (104). Qualitatively similar phase diagrams are expected when relaxing the unit-length constraint.
For $N_{f}=1$, the phase diagram is trivial: there is a single thermodynamic phase for any $N_{c}$ [41, 374, 43], without transitions. In this case we may only distinguish between different regimes: a confined regime for small values of $\beta J$ and $\beta\gamma$, and a Higgs regime for large $\beta J$ and $\beta\gamma$, which are analytically connected. For $\gamma=0$ it is easy to realize that no transitions are possible. Indeed, using the gauge and global symmetry, we can set (unitary gauge) $\Phi=(1,0,\ldots)$, obtaining a model of decoupled gauge fields. The same result holds for finite $\gamma$ [41].
For $N_{f}\geq 2$, as we have discussed in Sec. 8.1.2, there are various Higgs phases characterized by different spontaneous breakings of the global symmetry [99]. For $\beta\to\infty$, the symmetry of the ordered Higgs phases can be determined by looking at the minima of the scalar potential. Thus, results analogous to those presented in Sec. 8.1.2 are obtained. For $N_{f}\geq N_{c}$ there are two Higgs phases, one corresponding to negative values of $v$ for large $\beta$, and a second one corresponding to positive values of $v$. For $N_{f}<N_{c}$, there is instead a single ordered phase in the negative-$v$ region. Indeed, the large-$\beta$ region remains disordered for $v>0$.
@@ 9.1.2 Low-temperature Higgs phases
For $N_{f}\geq N_{c}$ a second Higgs phase occurs, since the system orders for $v>0$ when $\beta$ is increased. The nature of this phase depends on $N_{f}$ and $N_{c}$. For $N_{f}=N_{c}$ the global symmetry breaking pattern is $\mathrm{U}(N_{f})\rightarrow\mathrm{SU}(N_{f})$, cf. Eq. (89). In this case the SU($N_{f}$) symmetry is not broken in the Higgs phase, so the transition is only related to the breaking of the U(1) symmetry, the order parameter being $\mathrm{det}\,\Phi$, as discussed in Sec. 8.1.2. Thus, the system can develop continuous LGW transitions belonging to the XY universality class. No charged continuous transitions described by the corresponding NAHFT are expected, since charged FPs exist only for $N_{f}\gg N_{c}$ (unless some new features emerge in three dimensions, like the case of the one-component LAH model with noncompact gauge variables, see the discussion at the end of Sec. 5.2.2). Finally, for $N_{f}>N_{c}$ and $v>0$ we have transitions with symmetry-breaking pattern $\mathrm{U}(N_{f})\rightarrow\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$, cf. Eq. (90). In this case, charged continuous transitions corresponding to the stable FP of the corresponding NAHFT can be only observed for $N_{f}>N_{f}^{*}(3)$, where $N_{f}^{*}(3)$ is the ($N_{c}$-dependent) critical flavor number defined in Sec. 8.1.3.444444It is worth noting that a LGW field-theoretical description in terms of a gauge-invariant order-parameter field does not exist, since the Higgs phase symmetry $\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$ depends on the number of colors. Therefore, any effective description should somehow take into account some specific properties of the gauge interaction. In principle, this color-dependent symmetry breaking may be realized in LGW× transitions with a gauge-dependent order-parameter field. However, the latter scenario is quite unlikely in 3D systems, because gauge fluctuations are expected to be generally relevant at the uncharged FP of theories with continuous gauge groups, see Sec. 8.1.3.
@@ 9.2.1 Lattice SU(2) gauge Higgs models
We now focus on the specific case of the LNAH models with SU(2) gauge symmetry, which present some peculiar features that are not present for $N_{c}\geq 3$. The phase diagrams of multicomponent LNAH models with $N_{c}=2$ are sketched in Fig. 19 for a generic value of $\gamma$. For any $N_{f}\geq 2$ there are two low-temperature Higgs phases separated by a first-order transition line. Since the LNAH model (99) with $v=0$ is invariant under a larger group of transformations, the Sp($N_{f}$)$/\mathbb{Z}_{2}$ group [97, 167], the first-order line always lies on the line $v=0$. At the intersection point of the three transition lines, the behavior can be critical only for $N_{f}=2$, with a MCP belonging to the $\mathrm{Sp(}N_{2}\mathrm{)}/\mathbb{Z}_{2}=\mathrm{O(5)}$ universality class (see Sec. 9.2.2 for a more extensive discussion). For $N_{f}\geq 3$, RG arguments based on a multicritical LGW theory predict that the transitions at the intersection point are of first order [97, 167].
The behavior along the transition lines that separate the disordered phase from the ordered Higgs phases has been discussed in Sec. 9.1.2. For $v<0$, continuous transitions are only possible for $N_{f}=2$. They belong to the CP1, i.e., vector O(3), universality class (this has been numerically checked for $\gamma=0$ in Ref. [99]). For larger values of $N_{f}$, transitions are always of first order. For $v>0$, XY transitions are expected for $N_{f}=2$ (numerical results confirm this prediction [99]). For larger values of $N_{f}$, we may have charged continuous transitions associated with the stable charged FPs of the NAHFT, see Sec. 8.1.3. Numerical studies [99, 392] for relatively large values of $N_{f}$ confirm this scenario. Indeed, for a sufficiently large number $N_{f}$ of flavors the SU(2) gauge LNAH models undergo continuous transitions for finite (sufficiently large) values of $\gamma$, with a critical behavior consistent with that predicted by the 3D SU(2) gauge Higgs field theory, as confirmed by the nice agreement between numerical lattice results and large-$N_{f}$ computations, see also Sec. 9.3.

### v3

@@ 8.2 SO($N_{c}$) gauge Higgs theory with scalar fields in the fundamental representation
The NAHFT with gauge group SO($N_{c}$) and multiflavor scalar fields in the fundamental representation is obtained by considering a real scalar field $\Phi^{af}({\bm{x}})$ (with $a=1,...,N_{c}$ and $f=1,\ldots,N_{f}$) and the gauge field $A_{\mu}^{c}({\bm{x}})$, $c=1,\ldots,N_{c}(N_{c}-1)/2$. The most general renormalizable Lagrangian is [408, 304, 186]
${\cal L}_{\rm SO}={1\over 2g^{2}}{\rm Tr}\,{\cal F}_{\mu\nu}^{2}++{\rm Tr}\,[(D_{\mu}\Phi)^{\dagger}(D_{\mu}\Phi)]+V(\Phi),$ (97)
@@
where ${\cal F}_{\mu\nu}$ is the non-Abelian field strength of the gauge field ${\cal A}_{\mu}=\sum_{c}A_{\mu}^{c}T^{c}$, $D_{\mu}$ is the covariant derivative in the fundamental representation of SO($N_{c}$), and $T_{c}$ are the generators of the SO($N_{c}$) algebra. For any value of $N_{c}$ and $N_{f}$, the Lagrangian (97) is invariant under the local SO($N_{c}$) gauge transformations,
@@
The RG flow of the theory does not have stable charged FPs with $\alpha>0$, unless $N_{f}>N_{f}^{*}(d)$, where $N_{f}^{*}(d)$ is very large in four dimensions. Indeed, we find $N^{*}_{f}(4)=210.5$ for $N_{c}=2$, and $N^{*}_{f}(4)\approx 443.8$, for $N_{c}=3$. The stable FPs always lie in the region $u,v>0$. More precisely, one finds $v^{*}\approx 6\varepsilon/N_{f}$, $u^{*}\approx 6\varepsilon/N_{f}$, and $\alpha^{*}=12\varepsilon/N_{f}$ for large values of $N_{f}$. Note that the existence of a critical theory for large values of $N_{f}$ is also confirmed by the analysis of the model using the large-$N_{f}$ expansion [304].
The results reviewed above show that continuous transitions with the effective field-theory description (97) are only possible for a large number of components. Again, we remind the reader that this result does not exclude the presence of other types of continuous transitions (see Sec. 3.1) for small values of $N_{f}$.
@@ 8.3 SU($N_{c}$) gauge Higgs theory with scalar fields in the adjoint representation
To define the model we consider real scalar fields $\Phi^{af}({\bm{x}})$ with $a=1,...,N_{c}^{2}-1$ (color index) and $f=1,...,N_{f}$ (flavor index), and $N_{c}^{2}-1$ real gauge fields $A_{\mu}^{c}({\bm{x}})$. The most general renormalizable Lagrangian is
${\cal L}_{\rm SUadj}={1\over 2g^{2}}{\rm Tr}\,{\cal F}_{\mu\nu}^{2}+{\rm Tr}\,[(D_{\mu}\Phi)^{\dagger}(D_{\mu}\Phi)]+V(\Phi),$ (100)
@@
where ${\cal F}_{\mu\nu}$ is the field strength of the gauge field ${\cal A}_{\mu}\equiv\sum_{c}A_{\mu}^{c}T^{c}$, $D_{\mu}$ is the covariant derivative in the adjoint representation, i.e., $\partial_{\mu}+i\sum_{c}A_{\mu}^{c}T_{A}^{c}$, and $T_{A}^{abc}=-if^{abc}$ are the Hermitian SU($N_{c}$) generators in the adjoint representation ($f^{abc}$ are the structure constants of the SU($N_{c}$) group). The model is invariant under left SU($N_{c}$) local transformations and under right O($N_{f}$) global transformations.
@@ 9 Lattice non-Abelian Higgs models
As we shall see, LNAH models present various phases and phase transitions, whose nature depends of the global and local symmetries. As already discussed for LAH theories, see Secs. 5 and 6, one of the main issues is the identification of the effective field-theoretical description of the phase transitions present in LNAH models. As already remarked before, to define nonperturbatively the NAHFT as the continuum limit of a lattice model, it is crucial that the corresponding LNAH models, with the same local and global symmetries, undergo charged continuous transitions where both scalar and gauge modes are critical.
The phase transitions of lattice SU($N_{c}$) gauge Higgs models exhibit critical behaviors of various types, as discussed in Sec. 3.1. There are transitions that admit a LGW description in terms of a gauge-invariant order-parameter field—this is appropriate when gauge modes are not critical—and GFT transitions with critical gauge modes, associated with the stable charged FPs of the NAHFTs discussed in Sec. 8. We mostly focus on lattice models with scalar matter in the fundamental representation of the gauge group. But we also discuss the case of scalar fields that transform in the adjoint representation of the SU($N_{c}$) gauge group, where also topological transitions appear, which are not associated with the breaking of the global symmetry. Finally, we briefly mention the main features of the phase diagrams and phase transitions of LNAH models with SO($N_{c}$) gauge symmetries.
@@ 9.1 Lattice SU($N_{c}$) gauge Higgs models with multiflavor scalar fields
The Hamiltonian $H$ is the sum of the scalar-field kinetic term $H_{K}$, of the local scalar potential $H_{V}$, and of the pure-gauge Hamiltonian $H_{G}$. As usual, we set the lattice spacing equal to one, so all lengths are measured in units of the lattice spacing. The kinetic term $H_{K}$ is given by
@@
The Hamiltonian $H$ is invariant under global U($N_{f}$) transformations that only act on the scalar field, $\Phi\to\Phi W$, $W\in\hbox{U($N_{f}$)}$, and under local SU($N_{c}$) gauge transformations
$\Phi_{\bm{x}}\to V_{\bm{x}}\Phi_{\bm{x}},\qquad U_{{\bm{x}},\mu}\to V_{\bm{x}}U_{{\bm{x}},\mu}V_{{\bm{x}}+\hat{\mu}}^{\dagger},\qquad V_{\bm{x}}\in{\rm SU}(N_{c}).$ (106)
The lattice model (102) can be simplified by considering the fixed-length limit of the scalar field, i.e., by requiring ${\rm Tr}\,\Phi_{\bm{x}}^{\dagger}\Phi_{\bm{x}}^{\phantom{\dagger}}=1$, which can be formally obtained by considering the limit $u,r\to\infty$ keeping the ratio $r/u=-1$ fixed. In the unit-length limit the lattice potential (104) reduces to
@@
The model in the unit-length limit is generally expected to have the same features as models with generic $r$ and $u$.
@@
where $\epsilon^{ab}$ is the completely antisymmetric tensor in the color space, with $\epsilon^{12}=1$. Note that, if $N_{f}>N_{c}=2$, the composite operator $Y_{\bm{x}}^{fg}$ transforms nontrivially under the SU($N_{f}$) group, therefore the U(1) symmetry can be spontaneously broken only in phases characterized by a broken SU($N_{f}$) symmetry. Instead, for $N_{f}=N_{c}=2$, the operator $Y_{\bm{x}}^{fg}$ can be rewritten as $Y_{\bm{x}}^{fg}=\epsilon^{fg}\det\Phi_{\bm{x}}$, and is therefore invariant under SU(2) global transformations. Thus, the U(1) symmetry can be broken irrespectively of the breaking of the SU($N_{f}$) symmetry. Analogous results hold for generic models with $N_{f}=N_{c}$, irrespective of the value of $N_{c}$, see Refs. [198, 100].
We now outline the main features of the phase diagram of the LNAH models (102), discussing the nature of their phase transitions and critical behaviors. We mainly focus on the LNAH models satisfying the unit-length constraint (107). Qualitatively similar phase diagrams are expected when relaxing the unit-length constraint, i.e. for generic SU($N_{c}$) gauge LNAH models with Hamiltonian (102).
9.2.1 General considerations on the low-temperature Higgs phases
For $N_{f}=1$, the phase diagram of the SU($N_{c}$) gauge LNAH models is trivial: there is a single thermodynamic phase for any $N_{c}$ [41, 406, 43], without transitions. In this case we may only distinguish two different regimes: a confined regime for small values of $\beta J$ and $\beta\gamma$, and a Higgs regime for large $\beta J$ and $\beta\gamma$, which are analytically connected. For $\gamma=0$ it is easy to realize that no transitions are possible. Indeed, using the gauge and global symmetry, we can set (unitary gauge) $\Phi=(1,0,\ldots)$, obtaining a model of decoupled gauge fields. The same result holds for finite $\gamma$ [41].
For $N_{f}\geq 2$, as we have discussed in Sec. 8.1.2, there are various Higgs phases characterized by different spontaneous breakings of the global symmetry [100]. For $\beta\to\infty$, the symmetry of the ordered Higgs phases can be determined by looking at the minima of the scalar potential. Thus, results analogous to those presented in Sec. 8.1.2 are obtained. For $N_{f}\geq N_{c}$ there are two Higgs phases, one corresponding to negative values of $v$ for large $\beta$, and a second one corresponding to positive values of $v$. For $N_{f}<N_{c}$, there is instead a single ordered phase in the negative-$v$ region. Indeed, for large values of $\beta$ the system is always disordered for $v>0$.
@@ 9.2.1 General considerations on the low-temperature Higgs phases
For $N_{f}\geq N_{c}$ a second Higgs phase occurs, since the system orders for $v>0$ when $\beta$ is increased. The nature of this phase depends on $N_{f}$ and $N_{c}$. For $N_{f}=N_{c}$ the global symmetry breaking pattern is $\mathrm{U}(N_{f})\rightarrow\mathrm{SU}(N_{f})$, cf. Eq. (92). In this case the SU($N_{f}$) symmetry is not broken in the Higgs phase, so the transition is only related to the breaking of the U(1) symmetry, the order parameter being $\mathrm{det}\,\Phi$, as discussed in Sec. 8.1.2. Thus, the system can undergo continuous LGW transitions belonging to the XY universality class. No charged continuous transitions described by the corresponding NAHFT are expected, since charged FPs exist only for $N_{f}\gg N_{c}$ (unless some new features emerge in three dimensions, as it occurs for the one-component LAH model with noncompact gauge variables; see the discussion at the end of Sec. 5.2.2). Finally, for $N_{f}>N_{c}$ and $v>0$ we have transitions with symmetry-breaking pattern $\mathrm{U}(N_{f})\rightarrow\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$, cf. Eq. (93). In this case, charged continuous transitions corresponding to the stable FP of the corresponding NAHFT can be observed for $N_{f}>N_{f}^{*}(3)$, where $N_{f}^{*}(3)$ is the ($N_{c}$-dependent) critical flavor number defined in Sec. 8.1.3.464646It is worth noting that a LGW field-theoretical description in terms of a gauge-invariant order-parameter field does not exist, since the Higgs phase symmetry $\mathrm{SU}(N_{c})\otimes\mathrm{U}(N_{f}-N_{c})$ depends on the number of colors. Therefore, any effective description should somehow take into account some specific properties of the gauge interaction. In principle, this color-dependent symmetry breaking may be realized in LGW× transitions with a gauge-dependent order-parameter field. However, the latter scenario is quite unlikely in 3D systems, because gauge fluctuations are expected to be generally relevant at the uncharged FP of theories with continuous gauge groups, see Sec. 8.1.3.
@@ 9.2.2 Lattice SU(2) gauge Higgs models
We now focus on LNAH models with SU(2) gauge symmetry, which present some peculiar features that are not present for $N_{c}\geq 3$. The phase diagrams of multicomponent LNAH models with $N_{c}=2$ are sketched in Fig. 19 for a generic value of $\gamma$. For any $N_{f}\geq 2$ there are two low-temperature Higgs phases separated by a first-order transition line. Since the LNAH model (102) with $v=0$ is invariant under a larger group of transformations, the Sp($N_{f}$)$/\mathbb{Z}_{2}$ group [98, 198], the first-order line always lies on the line $v=0$. At the intersection point of the three transition lines, the behavior can be critical only for $N_{f}=2$, with a MCP belonging to the $\mathrm{Sp(}N_{2}\mathrm{)}/\mathbb{Z}_{2}=\mathrm{O(5)}$ universality class (see Sec. 9.2.3 for a more extensive discussion). For not too large values of $N_{f}$ satisfying $N_{f}\geq 3$, [^47] We should mention that the field-theory model with $v=0$ admits a charged FP for large values of $N_{f}$. Thus, for $N_{f}\gg 1$ a charged multicritical behavior can also appear. RG arguments based on a multicritical LGW theory predict the intersection point to correspond to a first-order transition [98, 198].
The behavior along the transition lines that separate the disordered phase from the ordered Higgs phases has been discussed in Sec. 9.2.1. For $v<0$, continuous transitions are only possible for $N_{f}=2$. They belong to the CP1, i.e., vector O(3), universality class (this has been numerically checked for $\gamma=0$ in Ref. [100]). For larger values of $N_{f}$, transitions are always of first order. For $v>0$, XY transitions are expected for $N_{f}=2$ (numerical results confirm this prediction [100]). For larger values of $N_{f}$, we may have charged continuous transitions associated with the stable charged FPs of the NAHFT, see Sec. 8.1.3. Numerical studies [100, 434] for relatively large values of $N_{f}$ confirm this scenario. Indeed, for a sufficiently large number $N_{f}$ of flavors the SU(2) gauge LNAH models undergo continuous transitions for finite (sufficiently large) values of $\gamma$, with a critical behavior consistent with that predicted by the 3D SU(2) gauge Higgs field theory. This is confirmed by the comparison of numerical lattice results and large-$N_{f}$ field-theoretical predictions, see also Sec. 9.3.